OpenStudy (anonymous):

which is greater 5 over 9 or 3 over 5?

OpenStudy (anonymous):

Which one do you think is greater?

OpenStudy (anonymous):

\[\frac{5}{9}<\frac{3}{5}\] 5/9=0.55... and 3/5=0.6